Part A: For the reaction 2 A + B → C, the initial rate is measured at several different reactant concentrations. From the resulting tabulated data, determine the rate constant for the reaction. [A](M) [B](M) Initial Rate (M/s) 0.05 0.05 0.035

0.10 0.05 0.070 0.20 0.10 0.28

Answers?: 1.4×102M−2⋅s−1,   2.8×102M−2⋅s−1, 28M−2⋅s−1, 14M−1⋅s−1 Which one of these is it, and how?

Part B: The aquation of tris(1,10-phenanthroline)iron(II) in acid solution takes place according to the equation:
Fe(phen)32+ + 3 H3O+ + 3 H2O → Fe(H2O)62+ + 3 phenH+.
If the activation energy, Ea, is 126 kJ/mol and the rate constant at 30°C is 9.8 × 10-3 min-1, what is the rate constant at 35°C?

Part C: What are the units of k in the following rate law?
Rate = k[X]?

I have been stuck on these for a little while now, Thanks a lot for the help!
Answer #1

part B

ln(k2/k1) = Ea/R[1/T1-1/T2]

ln(x/(9.8*10^-3)) = (126*10^3/8.314)((1/303)-(1/308))

x = k at 35c = 2.2*10^-2 min-1.


part C

Rate = k[X]

rate units = M/S

[X] = M

K units = s-1